OpenStudy (anonymous):

Licence plates consist of either 3 letters and 3 numbers or 4 letters and 3 numbers. How many different licence plates can be issued..

OpenStudy (anonymous):

now we have an addition because of the "OR"

OpenStudy (anonymous):

Trying this again using nCr : So... 26c3 + 10c3 = 3496 26c4 + 10c3 = 23751 + 220 = 23971 = 27467 ?

OpenStudy (anonymous):

lol this nCr business is killing you i see it does not say that you cannot repeat numbers or letters

OpenStudy (anonymous):

oh sheesh.... so I might have been right before....and yes it really is

OpenStudy (anonymous):

so.... 26x26x26x10x10x10 + 26x26x26x26x10x10x10 ?

OpenStudy (anonymous):

counting principle all the way 26 choices for each of 3 letters gives a total of \(26^3\) choices for the letters

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

474552000 possibilities..

OpenStudy (anonymous):

whatever that number is, your first answer is correct

OpenStudy (anonymous):

phew..okay so to clarify for the 23545th time... nCr is used when there can NOT be repetition

OpenStudy (anonymous):

@geekchloe don't be confused by "formulas" make sure to understand what exactly is the question, and apply the counting principle. it is the basis for all of these

OpenStudy (anonymous):

yes, in general you have nCr when there cannot be repetition, such as picking marbles out of a bag and not replacing them or dealing cards out of a deck (since you don't deal a card, put it back, and then deal another one)

OpenStudy (anonymous):

right...that part makes sense

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

unfortunately..i probably still have more questions. thanks :P

OpenStudy (anonymous):

np